A person in a kayak starts paddling, and it accelerates from 0 to 0.567 m/s in a distance of 0.326 m. If the combined mass of the person and the kayak is 74.2 kg, what is the magnitude of the net force acting on the kayak?
Answer #1

U = 0 , V = 0.567m/s , S = 0.326m

=) a = V^2 /2S = 0.567^2 /2×0.326 = 0.493m/s^2

Now F = ma = 74.2×0.493 = 36.58N